# Large Multimodal Models for Low-Resource Languages: A Survey

### v1

@@ 4 Synthetic Data Generation
Another approach to efficiently create multimodal datasets for under-represented languages relies on synthetic data generation. While the previous section captures efforts comprising significant human involvement, we next cover synthetic data generation, which relies on existing resources and automated techniques to generate new data.
Figure 4: An overview of various fusion strategies employed in LMMs, categorized into early fusion, late fusion, and architectural fusion approaches. Early fusion combines features from different modalities (text, audio, and visual) using feature extractors and fusion techniques, before passing them to a classifier for the final output. Concatenation fusion directly concatenates features from different modalities, while gated fusion employs a gate controller network to regulate information flow between modalities. Late fusion processes each modality using separate models, then combines their predictions using decision-level fusion methods, such majority voting or weighted averaging. Architectural fusion approaches, such as attention fusion and encoder-decoder fusion, provide more sophisticated methods for multimodal integration. Attention fusion leverages self-attention layers and learned attention weights to selectively focus on relevant features across modalities.
@@
Back-translation. A common approach for synthetic data generation relies on the usage of back-translation, which has proven to be an effective technique to enhance the data for multilingual MT (MMT) in LR language pairs. Chowdhury et al. [2018] demonstrated the effectiveness of this technique for training a neural MMT system in the context of LR language pairs by leveraging the Flickr30k dataset Young et al. (2014) and translating the source language (English) captions to the target LR language (Hindi). In the WMT24 English-to-Low-Resource Multi-Modal Translation task, Haq et al. [2024] showcased the effectiveness of back-translation for Hindi. Another use case of back-translation was shown by Alwajih et al. [2024], who, starting from English-based image-text pairs, employed translation to Arabic, as well as back-translation. This was necessary for evaluating the quality of the translation, before passing the data to humans for Arabic dialect translation and training a dialect-aware LMM, named Dallah. However, the consistency of back-translated synthetic data has been a concern. To address this issue, Wang et al. [2023b] proposed a framework to improve the robustness of models when adapting grounded VQA models to LR languages, aiming to improve the performance without relying on machine-translated data.
@@
Image-based generation. Another mainstream approach for synthetic data generation uses images as a starting point. In the case of Indic language multimodal MT Dash et al. (2023), synthetic images generated by diffusion models were deemed beneficial, their main goal being that of capturing the complexity of the target domain, and augmenting the existing image dataset. Similarly, Haq et al. [2024] created exhaustive image descriptions in addition to the already existing short region-based descriptions. Doan et al. [2024] utilized image-based generation for several purposes, such as description generation and relevant information extraction, to develop Vintern-1B, an efficient LMM for Vietnamese. Collectively, these studies demonstrate the versatility and effectiveness of synthetic data for tackling a diverse set of multimodal tasks in the context of LR languages.
@@ 5 Multimodal Fusion Techniques
We identified three distinct types of fusion approaches employed in multimodal learning, categorized into early fusion, late fusion, and architectural fusion approaches. An overview of the different fusion strategies is provided in Figure 4. The diagram depicts the various ways in which textual, visual and auditory features can be combined at different stages to enable effective integration of multimodal information.
Early fusion. Early fusion, also known as feature-level fusion, involves combining features from different modalities at the input level before passing them through a unified model. In Persian sentiment analysis, Dashtipour et al. [2021] demonstrated the effectiveness of early fusion by combining acoustic, visual, and textual features through a context-aware framework. Their approach showed that integrating features at an early stage allowed the model to capture cross-modal interactions more effectively, achieving 91.39% accuracy compared to unimodal approaches. Similarly, the shared task on Tamil and Malayalam multimodal sentiment analysis Premjith et al. (2023) revealed that early fusion techniques were particularly effective for handling code-mixed content and cultural nuances specific to these languages.
@@
For Amharic hate speech detection in memes Jigar et al. (2024), the authors employed concatenation, directly combining visual features from memes with textual features, demonstrating the effectiveness of this straightforward approach for LR languages. The integration of multimodal features through gating mechanisms has shown particular promise in LR scenarios, as demonstrated in English-to-Low-Resource translation tasks for Hindi, Malayalam, Bengali, and Hausa, where Hatami et al. (2024) used gated fusion to selectively combine visual and textual information. This approach was further validated by Alalem et al. [2023] in their Audio-Text Fusion (ATFusion) model for English and Egyptian Arabic, where they employed Group Gated Fusion (GGF) to dynamically control the flow of information between modalities, achieving superior performance over traditional fusion methods with accuracy improvements of up to 76.21% for English and 70.79% for Egyptian Arabic.
Late fusion. Late fusion, also known as decision-level fusion, combines predictions from separate modality-specific models at the decision stage rather than fusing features early in the pipeline. In this approach, individual models are trained independently for each modality (e.g., text, audio, and video) and their outputs are merged to produce the final prediction Zhang et al. (2024); Arifin et al. (2024). This allows each modality to be processed optimally according to its characteristics before integration. The independence between modality-specific models also provides flexibility, as individual components can be modified without affecting the others, and the system can still function if one modality is missing.
@@
Two popular late fusion strategies are weighted averaging and majority voting. In weighted averaging fusion, the predictions from different modalities are combined using weighted sums, with weights determining the contribution of each modality to the final decision Rahman et al. (2024); Karim et al. (2022). The weights can be uniform or learned to optimize performance. Majority voting fusion employs gating mechanisms to control information flow between modalities and determine which modality should be emphasized. For example, Dashtipour et al. [2021] used gating networks to adaptively combine predictions from audio, visual and textual models based on their estimated reliability for Persian sentiment analysis. Their results showed that intelligent fusion using gates improved performance compared to simple averaging, highlighting the benefits of selective information integration in multimodal systems.
@@
Architectural fusion. Encoder-decoder fusion architectures have emerged as an effective approach for multimodal integration across different languages and tasks. Several studies have demonstrated success with this approach. For example, Chakravarthi et al. [2019] employed an encoder-decoder framework with phonetic transcription to improve machine translation between Dravidian languages, while Sehar et al. [2021] utilized an encoder-decoder architecture to fuse audio, video and text features for Urdu sentiment analysis. Similarly, Meetei et al. [2024] showed that encoder-decoder fusion of correlated modalities can enhance translation quality for LR languages. The key advantage of encoder-decoder architectures is their ability to first encode input features from different modalities into a shared representation space before decoding them into the target output.
@@
Attention-based fusion has also proven to be highly effective for multimodal integration Ristea and Ionescu (2023). As shown by Haputhanthri et al. [2023] for Sinhala sign language recognition, attention mechanisms allow the model to dynamically focus on the most relevant features across modalities. Yang et al. [2024] successfully employed attention fusion for Mongolian sentiment analysis by combining features from audio, text and visual inputs. Zhang et al. [2024] demonstrated that attention-based fusion of multimodal data improves depression risk detection by allowing the model to attend to salient information across audio, video and text modalities. The ability of attention mechanisms to learn dynamic weights between modalities makes them particularly suitable for tasks requiring adaptive integration of complementary information sources.
@@
Comparative analysis of fusion techniques. Each fusion approach presents distinct advantages and challenges in the context of LR languages. Early fusion enables deep interaction between modalities from the start, but can be computationally expensive and may suffer when one modality is noisy. Late fusion offers flexibility and robustness when modalities are missing, but may miss important cross-modal interactions. Architectural fusion approaches show promise in capturing complex relationships between modalities, but require careful tuning and substantial computational resources. A notable innovation in this space is the Multi-Representative Fusion (MRF) mechanism Chauhan et al. (2022), which generates diverse representations for each modality and selectively chooses the best fusion via attention. This approach has shown particular promise in handling noisy inputs, achieving state-of-the-art performance on several LR sentiment analysis benchmarks.

### v2

@@ 4 Synthetic Data Generation
An alternative approach to efficiently create multimodal datasets for LR languages relies on synthetic data generation. Unlike traditional dataset creation, which typically involves intensive manual data collection, human annotation, and domain-specific curation, synthetic data generation leverages existing resources and automated techniques to produce new multimodal content with minimal human input. This distinction is critical, as synthetic methods offer a scalable alternative for low-resource settings, where manual annotation is often costly or infeasible.
@@
Back-translation. A common approach for synthetic data generation relies on the usage of back-translation, which has proven to be an effective technique to enhance the data for multilingual MT (MMT) in LR language pairs. This technique works by translating text from an HR language into an LR language, and then back again, helping to generate additional aligned examples without requiring human involvement. Dutta Chowdhury et al. [2018-dutta] demonstrated the effectiveness of this technique for training a neural MMT system in the context of LR language pairs by leveraging the Flickr30k dataset [young2014image] and translating the source-language (English) captions to the target LR language (Hindi). Meetei et al. [meetei2021low] extended this approach for low-resource multimodal neural machine translation in the news domain for English-Hindi. In the WMT24 English-to-Low-Resource Multi-Modal Translation task, Haq et al. [2024-haq] showcased the effectiveness of back-translation for Hindi. Another use case of back-translation was shown by Alwajih et al. [2024-Alwajih-dallah], who, starting from English-based image-text pairs, employed translation to Arabic, as well as back-translation. This was necessary for evaluating the quality of the translation, before passing the data to humans for Arabic dialect translation and training a dialect-aware LMM, named Dallah. However, the consistency of back-translated synthetic data has been a concern. To address this issue, Wang et al. [2023-wang-adapting] proposed a framework to improve the robustness of models when adapting grounded VQA models to LR languages, aiming to improve the performance without relying on machine-translated data. Wang et al. [wang2022cross] further explored this challenge by introducing noise-robust learning for cross-lingual cross-modal retrieval to handle translation noise in machine-translated sentences.
@@
Image-based generation. Another mainstream approach for synthetic data generation uses images as a starting point. In the case of Indic language multimodal MT [2023-dash], synthetic images generated by diffusion models were deemed beneficial, their main goal being that of capturing the complexity of the target domain, and augmenting the existing image dataset. Similarly, Haq et al. [2024-haq] created exhaustive image descriptions in addition to the already existing short region-based descriptions. Doan et al. [2024-doan] utilized image-based generation for several purposes, such as description generation and relevant information extraction, to develop Vintern-1B, an efficient LMM for Vietnamese. Nath et al. [nath2022image] applied this approach for image caption generation in the low-resource Assamese language using an encoder-decoder framework that combines CNNs and RNNs to generate descriptions from images. Jiang et al. [jiang2024multimodal] expanded these approaches with multimodal seed data augmentation for the low-resource audio Latin Cuengh language, demonstrating how seed data can enhance intelligent recognition and comprehension of low-resource dialects. Collectively, these studies demonstrate the versatility and effectiveness of synthetic data for tackling a diverse set of multimodal tasks in the context of LR languages.
An emerging approach that avoids both traditional back-translation or image-based methods consists in leveraging the outputs of Vision-Language Models (VLMs). Qu et al. [2024-qu] generated multilingual responses for image-text inputs, translated them into English, and compared them with trusted references to detect hallucinations. These automatically mined hallucination-aware pairs are then used for direct preference optimization [2023-Rafailov-NeurIPS], enabling scalable fine-tuning without manual annotations, especially useful for low-resource languages.
Another innovative approach focuses on optimizing the composition of training data itself. Shukor et al. [shukor2025scaling_data] developed systematic methods to determine optimal domain weights for multimodal pre-training using scaling laws, validating their approach across Large Language Models, Native Multimodal Models, and Large Vision Models. This methodology provides principled alternatives to costly trial-and-error approaches for data mixture optimization, particularly valuable in resource-constrained settings, typical for low-resource language development.
Data sovereignty concerns. Synthetic data generation introduces risks beyond technical quality. Back-translation and LLM-based augmentation propagate source-language biases to target languages, potentially encoding cultural assumptions misaligned with target communities [navigli2023biases]. More critically, the CARE principles [carroll2020care] assert that Indigenous communities must retain authority over their linguistic data, a requirement that synthetic generation pipelines rarely accommodate. Evidence from Sámi language technology demonstrates the consequences: LLMs trained on available corpora without community oversight produce outputs that appear valid to non-speakers, but constitute nonsense to native speakers [wiechetek2024ethical]. We thus recommend that synthetic data pipelines incorporate community validation protocols and explicit data governance agreements prior to deployment.
Figure 5: An overview of various fusion strategies employed in LMMs, categorized into early fusion, late fusion, and architectural fusion approaches. Early fusion combines features from different modalities (text, audio, and visual) using feature extractors and fusion techniques, before passing them to a classifier for the final output. Concatenation fusion directly concatenates features from different modalities, while gated fusion employs a gate controller network to regulate information flow between modalities. Late fusion processes each modality using separate models, then combines their predictions using decision-level fusion methods, such as majority voting or weighted averaging. Architectural fusion approaches, such as attention fusion and encoder-decoder fusion, provide more sophisticated methods for multimodal integration. Attention fusion leverages self-attention layers and learned attention weights to selectively focus on relevant features across modalities.
@@ 5 Multimodal Fusion Techniques
Multimodal fusion refers to the process of combining information from different modalities (such as text, images, audio) to make more informed predictions or generate better outputs. Fusion can be seen as the “meeting point” where information from separate sensory channels comes together, similar to how humans integrate what they see, hear and smell to understand their environment. The choice of fusion strategy significantly impacts performance, especially in low-resource settings, where each modality might provide crucial complementary information that others lack. Below, we describe the primary approaches to fusion that represent different philosophies about when and how this integration should occur.
@@
We identify three distinct types of fusion approaches employed in multimodal learning, categorized into early fusion, late fusion, and architectural fusion approaches. An overview of the different fusion strategies is provided in Figure 5. The diagram depicts the various ways in which textual, visual and auditory features can be combined at different stages to enable effective integration of multimodal information. In Table 3, we provide a summary of computational requirements and efficiency trade-offs for a series of representative fusion approaches. We further discuss each fusion strategy independently, referring to the computational requirements and trade-offs along the way.
Early fusion. Early fusion, also known as feature-level fusion, involves combining features from different modalities at the input level before passing them through a unified model [youcef2021arabic, 2024-roken]. Early fusion can be conceptualized as “combining ingredients before cooking”, i.e. all modalities are mixed at the beginning of the processing pipeline. This allows the model to learn cross-modal interactions from the start, potentially capturing subtle relationships between modalities.
In Persian sentiment analysis, Dashtipour et al. [2021-dashtipour] demonstrated the effectiveness of early fusion by combining acoustic, visual, and textual features through a context-aware framework, achieving 91.39% accuracy with A+V+T concatenation. Similarly, Al-Azani et al. [2020-azani] showed that early fusion of textual, auditory, and visual modalities achieved over 94% accuracy for Arabic sentiment analysis.
The shared task on Tamil and Malayalam multimodal sentiment analysis [2023-premjith] also revealed that early fusion techniques are particularly effective for handling code-mixed content and cultural nuances specific to these languages [kodali2025bytesizedllm].
@@
For Amharic hate speech detection in memes, Jigar et al. [2024-Jigar] employed concatenation, directly combining visual features from memes with textual features, achieving 75% accuracy and demonstrating the effectiveness of this straightforward approach for LR languages [2022-debele]. The integration of multimodal features through gating mechanisms has shown particular promise in LR scenarios, as demonstrated in English-to-Low-Resource translation tasks for Hindi, Malayalam, Bengali, and Hausa, where Hatami et al. [2024-Hatami] used gated fusion to selectively combine visual and textual information. This approach was further validated by Alalem et al. [2023-alalem] in their Audio-Text Fusion model for English and Egyptian Arabic, where they employed Group Gated Fusion to dynamically control the flow of information between modalities, achieving superior performance over traditional fusion methods.
[…]
Table 3: Computational requirements and efficiency trade-offs for multimodal fusion techniques in low-resource settings. Bold values indicate configurations accessible for researchers with limited computational resources. A dash line indicates that the respective information is not specified in the original publication.
Method/Model Fusion Type GPU Req. Training Params Performance Key Trade-off Early Fusion Approaches MRF [2022-chauhan] Early 1080Ti 11GB 100 ep. $\approx$50M 84.1% Acc Noise-robust; needs multiple representations ViT + mBERT [faria2025sentimentformer] Early 40 ep. $\approx$200M 72.4% Acc High #params; moderate accuracy Swin + XLM-RoBERTa [faria2025sentimentformer] Early 40 ep. $\approx$280M 75.8% Acc Best early fusion; heavier A+V+T Concat [2021-dashtipour] Early $\approx$30M 91.4% Acc Simple; sync-sensitive BiLSTM Multimodal [2024-Jigar] Early (Concat) 32-64 ep. $\approx$10M 75.0% Acc Low #params; limited complexity Late Fusion Approaches XLM-R + DenseNet [2022-karim] Late GTX 1050 5-fold CV $\approx$400M 83.0% F1 Best multimodal; high memory MARBERTv2 + Ensemble [2023-albalawi] Late 100 ep. $\approx$180M 85.6% Acc Robust to missing modalities Intermediate / Architectural Fusion SentimentFormer [faria2025sentimentformer] Intermediate 30 ep. $\approx$220M 79.0% Acc Best overall; balanced cost AVTF-TBN [zhang-2024] Attention RTX 3090 24GB 300 ep. $\approx$100M 78.0% F1 High compute; medium accuracy CNN-LSTM Tagalog [deocampo2024lip] Intermediate 12 h $\approx$20M 89.5% Acc 25% faster than A+V Encoder–Decoder & Advanced Fusion URSA (3D-CNN + BLSTM) [2021-sehar] Feature-level 128+64 cells 95.4% Acc Feature >> decision fusion Feature-Extract [2024-arifin] Sep.+Merge T4 16GB 8.48M 93.3% Acc Low #params; specialized pipeline
Late fusion. Late fusion, also known as decision-level fusion, combines predictions from separate modality-specific models at the decision stage rather than fusing features early in the pipeline [deocampo2024lip, mamyrbayev2020multimodal]. Late fusion can be conceptualized as “requesting multiple expert opinions and then voting on a final decision” [zhang-2024, 2024-arifin]. In this approach, each modality is processed by its own specialized model, which becomes an expert in that particular type of data. Only after these individual experts have made their predictions are the results combined. This is particularly valuable when certain modalities might be missing or corrupted in real-world applications [2023-elahi].
@@
Two popular late fusion strategies are weighted averaging and majority voting. In weighted averaging, the predictions from different modalities are combined using a weighted sum, with weights determining the contribution of each modality to the final decision [2024-rahman, 2022-karim]. The weights can be uniform or learned to optimize performance. Majority voting employs gating mechanisms to control information flow between modalities and determine which modality should be emphasized [das2022multi]. For example, Dashtipour et al. [2021-dashtipour] used gating networks to adaptively combine predictions from audio, visual and textual models based on their estimated reliability for Persian sentiment analysis. Their results showed that intelligent fusion using gates improved performance compared with simple averaging, highlighting the benefits of selective information integration in multimodal systems.
Late fusion strategies are especially suitable for resource-constrained environments due to their flexibility and lower memory requirements. Since each modality is processed by independent models, the system can continue functioning when one modality is unavailable, enabling graceful degradation with missing inputs [2023-elahi, mamyrbayev2020multimodal]. For Arabic rumor detection, Albalawi et al. [2023-albalawi] achieved 83.83% accuracy with late fusion (MARBERTv2 + VGG-19 ensemble), compared with 85.57% for early fusion, demonstrating that the 1.7% performance gap is often smaller than the computational cost savings. Late fusion also enables parallel training of modality-specific models, reducing wall-clock time by 25-40% compared with end-to-end early fusion training [deocampo2024lip].
@@
Architectural fusion. Architectural fusion comprises more sophisticated integration methods that go beyond simple concatenation or averaging of features. Encoder-decoder fusion can be understood as a “translation system” between modalities, i.e. information from each modality is first converted into a common “language” (shared representation space) by encoders, before being decoded into the final output [faria2025sentimentformer, 2024-roken]. This allows the model to find complex mappings between very different data types. For example, Chakravarthi et al. [2019-chakravarthi] employed an encoder-decoder framework with phonetic transcription to improve machine translation between Dravidian languages, while Sehar et al. [2021-sehar] utilized an encoder-decoder architecture to fuse audio, video and text features for Urdu sentiment analysis. Similarly, Meetei et al. [2023-meetei] showed that encoder-decoder fusion of correlated modalities can enhance translation quality for LR languages. The key advantage of encoder-decoder architectures is their ability to first encode input features from different modalities into a shared representation space before decoding them into the target output.
@@
Attention-based fusion has also proven to be highly effective for multimodal integration [ristea2023cascaded]. As shown by Haputhanthri et al. [haputhanthri-2023] for Sinhala sign language recognition, attention mechanisms allow the model to dynamically focus on the most relevant features across modalities. Yang et al. [2024-yang] successfully employed attention fusion for Mongolian sentiment analysis by combining features from audio, text and visual inputs. Zhang et al. [zhang-2024] demonstrated that attention-based fusion of multimodal data improves depression risk detection by allowing the model to attend to salient information across audio, video and text modalities. The ability of attention mechanisms to learn dynamic weights between modalities makes them particularly suitable for tasks requiring adaptive integration of complementary sources.
Intermediate and architectural fusion approaches offer a balance between performance and accessibility. SentimentFormer [faria2025sentimentformer] achieves the highest Bangla meme accuracy (79.04%) with only 30 epochs, outperforming both early (75.83%) and late fusion (74.80%) on the same dataset. At the high-resource end, attention-based models such as AVTF-TBN [zhang-2024] require an RTX 3090 (24 GB) and 300 epochs for clinical-grade depression detection accuracy.
@@
Comparative analysis of fusion techniques. Each fusion approach presents distinct advantages and challenges in the context of LR languages [youcef2021arabic, deocampo2024lip]. Early fusion enables deep interaction between modalities from the start, but can be computationally expensive and may suffer when one modality is noisy. Late fusion offers flexibility and robustness when modalities are missing, but may miss important cross-modal interactions [mamyrbayev2020multimodal]. Architectural fusion approaches show promise in capturing complex relationships between modalities, but require careful tuning and substantial computational resources. A notable innovation in this space is the Multi-Representative Fusion (MRF) mechanism [2022-chauhan], which generates diverse representations for each modality and selectively chooses the best fusion via attention. This approach has shown particular promise in handling noisy inputs, achieving state-of-the-art performance on several LR sentiment analysis benchmarks.
Handling noisy modalities. A critical consideration for real-world deployment is robustness to corrupted modalities. The MRF mechanism [2022-chauhan] addresses this by generating multiple diverse representations for each modality and using attention to select the most informative fusion. When acoustic features are corrupted, MRF automatically reduces their contribution (from approximately 15% to <<5% of the final prediction), maintaining robust performance. However, MRF fails when all three modalities are simultaneously noisy for utterances critical to prediction. For Javanese emotion recognition [2024-arifin], separate modality processing achieves an accuracy of 93.32% compared with 71.15% for joint processing, specifically because independent processing minimizes interference when one channel contains noise.
Architectural complexity considerations. Our analysis suggests that the additional complexity of architectural fusion is justified in three cases: (1) when cross-modal interactions are semantically rich and task-critical, as in Sinhala sign language recognition [haputhanthri-2023] and Mongolian sentiment analysis [2024-yang], where temporal alignment between visual gestures and linguistic features requires learned attention weights; (2) when modalities have different noise characteristics or information densities, as demonstrated for Urdu sentiment analysis where feature-level fusion (95.35%) substantially outperformed decision-level fusion (91.23%) [2021-sehar]; and (3) for clinical or safety-critical applications where prediction errors have serious consequences [zhang-2024]. Conversely, for rapid prototyping or tasks where text modality dominates (e.g. in Bengali hate speech detection, where a text-only XLM-RoBERTa achieves $F1=0.82$ vs. $F1=0.83$ for the multimodal pipeline [2022-karim]), simpler approaches may be preferable.
Table 4: Performance comparison of early, late and intermediate fusion for low-resource languages. Best score on each row is highlighted in bold.
Language/Task Early Late Intermediate Best Strategy Bangla Memes [faria2025sentimentformer] 75.83% 74.80% 79.04% Intermediate Arabic Rumors [2023-albalawi] 85.57% 83.83% Early Urdu Sentiment [2021-sehar] 95.35% 91.23% Early Javanese Emotion [2024-arifin] 71.15% 93.32% Separate Processing Amharic Memes [2024-Jigar] 75.00% Early Persian Video [2021-dashtipour] 91.39% 90.32% Early
Table 5: Decision guide for selecting the fusion strategy based on constraints and requirements. ✓✓ = Strongly recommended, ✓ = Suitable, $\sim$ = Acceptable, ✗ = Not recommended. Based on empirical findings from [2022-chauhan, faria2025sentimentformer, zhang-2024, mamyrbayev2020multimodal].
Requirement/Constraint Early Late Architectural
Missing modality robustness ✓✓
Noisy input handling ✓✓ (MRF)
Low memory (<<8GB VRAM) ✓✓ $\sim$
Fast training (<<50 epochs) $\sim$ ✓✓ $\sim$
Maximum accuracy ✓✓ $\sim$ ✓✓
Cross-modal interactions ✓✓ ✓✓
Rapid prototyping $\sim$ ✓✓
Clinical/safety-critical $\sim$ ✓✓
In Table 4 we present fusion strategy performance across different low-resource languages and tasks, while in Table 5, we provide a decision guide based on specific constraints. Early fusion generally achieves the highest accuracy (e.g. 95.35% for Urdu, 91.39% for Persian video analytics), but the performance gap between strategies is often smaller than the computational cost difference. For researchers with limited computational resources (single GPU, <<16GB VRAM), we recommend starting with lightweight early fusion models such as BiLSTM ($\approx$10M parameters) to establish baselines, before progressing to intermediate fusion with efficient architectures for improved performance.